# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ 7.1.7 Evaluation Metric for Text Datasets
  • Partial Match (PM) metric evaluates the similarity between a model’s output and the reference by allowing partial credit for partially correct answers. Unlike strict metrics like Exact Match (EM), PM accounts for overlaps or shared elements, such as keywords or phrases, making it more flexible in assessing performance.
@@
  • Recall is a metric used to evaluate a model’s ability to retrieve all relevant instances in a dataset. It is calculated as the ratio of correctly retrieved relevant items to the total number of relevant items, emphasizing completeness.
@@
  • F1 is a performance measure that combines Precision and Recall into a single score using their harmonic mean. It provides a balanced evaluation, especially useful in datasets with imbalanced classes, by considering both false positives and false negatives.
  • BLEU [288], is a widely used metric for evaluating the quality of machine-generated text, especially in machine translation. It works by comparing n-grams in the generated output with reference texts to measure overlap, while applying penalties for overly short outputs to ensure fluency.
  • SacreBLEU [289] is a standardized version of the BLEU metric used to evaluate machine translation quality. It simplifies BLEU’s implementation by fixing preprocessing steps like reference handling to ensure consistent and reproducible results across different systems.
  • Rouge [290] and its variants measure model’s performance by calculating overlap between model output and reference answer with unigram(Rouge-1), bigram(Rouge-2), LCS(Rouge-L), etc. Gold Rouge-1 in VCSUM dataset refers to using high-quality reference summaries (gold standards) for evaluation, ensuring reliable and meaningful comparisons.
@@
  • Edit Similarity is a metric that measures the similarity between two text sequences based on the minimum number of edit operations required to transform one sequence into another. It is derived from the concept of edit distance such as Levenshtein distance.
  • Pass@k [293] evaluates the performance of a model by measuring the percentage that at least one of the top k generated outputs contains a correct solution. In datasets we surveyed, only Pass@1 is used.
  • Exponential Similarity is a metric that measures the similarity between two items by exponentially weighting their differences, giving more importance to smaller discrepancies.
@@
  • Mean Reciprocal Rank (MRR) is an evaluation metric commonly used in information retrieval and recommendation systems to measure the quality of ranked results. It calculates the reciprocal of the rank of the first relevant item in a result list and averages it across all queries.
TABLE XIX: Multimodal Dataset. Specfically, for data type, Img: Image; T: text; V: Video. For task abbreviation, Conv: conversation task; Desc: description task; Reas: reasoning task; Perc: perception task; Pred: prediction; NTH: needle in the haystack; SUMM: summary. For instance and average column, Q: questions; W: words; s: seconds. For example, 54 Img, 150 Q denote that there are 54 images with 150 questions.
Tasks Name Data Source Instance Average Metric Language
Conv, Desc, Reas LLaVA-Bench [294] Img, T COCO, In-The-Wild 54 Img, 150 Q 1 Img, 59.9 W Relative Score EN
Perc, Reas MMBench [295] Img, T Internet 2948 Q 1 Img, 114.5 W Acc EN/CN
Pred, Count,
NIH, Retrieval
MileBench [296] Img, T
Public,
self-building
6440 Q 15.2 Img, 422.3 W Acc, ROUGE-L EN
Reas, NIH, SUMM,
Desc, Order, Count
MLVU [297] V, T
Public,
self-collection
1334 V, 2593 Q 704.6s V, 39.7 W M-Avg, G-Avg EN
Reas, Retrieval LongVideoBench [298] V, T web-collected 3763 V, 6678 Q 730.5s V, 49.5 W Acc EN
Perc, Recognition, Reas Video-MME [299] V, T YouTube 900 V, 2700 Q 1017.9s V Acc EN
Desc, Reas NExT-QA [300] V, T
YouTube,
TV Show, Public
1000 V, 47962 Q 44s V, 25.5 W Acc, WUPS EN
Perc, Count, Reas MVBench [301] V, T Public 4000 Q 16.7s V, 31.3 W Acc EN
Decs MSVD-QA [302] V, T MSVD 1970 V, 50505 Q 10s V Acc EN
Desc MSRVYY-QA [302] V, T MSRVTT 10000 V, 243690 Q 15s V Acc EN
7.2 Multimodal Datasets and Evaluation Metric
7.2.1 Multimodal Datasets
Multimodal datasets have emerged to address the need for a comprehensive understanding of the complex real world by integrating diverse data types such as text, images, audio, and video. These datasets drive advancements in AI, particularly in machine learning and deep learning, by offering rich and diverse data to train more robust and versatile models. We analyze the multimodal benchmarks listed in Table XIX, highlighting their distinct focuses. Each benchmark is built upon one or more multimodal datasets, involving their collection, processing, and the use of specific validation metrics. Below, we provide a detailed introduction and description of each multimodal benchmark.
  • LLaVA-Bench [294] The benchmark is structured around image-ground-truth textual description-question-answer triplets, segmented across COCO and In-The-Wild datasets. It assesses a model’s proficiency in multimodal instruction adherence and visual reasoning. By employing a suite of tasks and metrics, it quantifies the model’s ability to comprehend and act on visual-language directives, articulate comprehensive descriptions, and engage in intricate reasoning processes.
  • MMBench [295] This benchmark serves as a bilingual multimodal benchmark, facilitating a comparative analysis of VLM performance across English and Chinese linguistic contexts. It distinctively assesses multimodal models using a hierarchical taxonomy of abilities, stringent quality assurance measures, and a dual-language evaluation framework. Unlike other benchmarks, MMBench [295] incorporates the CircularEval strategy for comprehensive evaluation and utilizes LLMs for precise extraction of choices, setting it apart from its counterparts.
  • MileBench [296] evaluates the multi-modal long-context capabilities of LLMs, including both diagnostic and realistic evaluation sets. It emphasizes long-context and multi-image tasks. This unique focus allows it to capture the complexity and diversity of real-world multimodal challenges, setting it apart from existing benchmarks. The dataset in MileBench [296] is characterized by its inclusion of long texts integrated with multiple images, reflecting real-world scenarios where context is key. It contains a diverse range of tasks that require both comprehension and generation.
  • MLVU [297] is a holistic benchmark, designed to gauge the capabilities of multi-modal LLMs in comprehending video content, transcends the constraints of its predecessors by significantly increasing video durations, encompassing diverse video genres, and crafting a spectrum of assessment tasks. This benchmark offers an extensive array of tasks and video genres to evaluate the comprehensive competencies of MLLMs. It highlights the substantial potential for enhancement in current methodologies and emphasizes the critical factors of context length, image comprehension quality, and the selection of LLM architecture for future progress.
  • LongVideoBench [298] This benchmark offers an extensive benchmarking framework aimed at assessing the capacity of large multimodal models (LMMs) to comprehend lengthy videos with subtitles, extending up to an hour. It places a strong focus on the retrieval and reasoning capabilities over extended, interwoven video and language data streams, tackling the challenge of single-frame bias and underscoring its proficiency in evaluating multimodal comprehension in long contexts.
  • Video-MME [299] A benchmark for comprehensive evaluation, it assesses the proficiency of Multi-modal Large Language Models (MLLMs) in analyzing videos. This dataset comprises a wide array of 900 videos spanning diverse domains and subfields, ensuring extensive scenario coverage. It encompasses videos with lengths ranging from 11 seconds to 1 hour to gauge model flexibility across various time frames. Furthermore, it incorporates various data modalities, including subtitles and audio tracks, to evaluate the comprehensive competencies of MLLMs. The benchmark aims to test the models’ capacity for sequential visual data comprehension, with an emphasis on temporal reasoning and the processing of multi-modal inputs.
  • NExT-QA [300] Advancing video comprehension from mere description to explanation of causal, temporal, and descriptive actions, a video question answering (VideoQA) benchmark has been established. This benchmark boasts a dataset with 5,440 videos and approximately 52K manually annotated question-answer pairs, sorted into causal, temporal, and descriptive categories. It poses a challenge to QA models to engage in reasoning about causal and temporal actions and to decipher complex object interactions within daily activities. Distinguished from other video benchmarks, this benchmark specifically focuses on causal and temporal action reasoning within realistic videos that are rich in object interactions. It stands as one of the largest manually annotated VideoQA datasets, offering support for both multiple-choice and open-ended questions, and includes a variety of videos that mirror real-life scenarios.
  • MVBench [301] Featuring a substantial dataset, the benchmark comprises 200 multiple-choice question-answer (QA) pairs for each of the 20 temporal understanding tasks, amassing a total of 4,000 QA pairs. It draws from a variety of videos across 11 public datasets, spanning diverse domains and scenes, thereby testing models’ abilities to comprehend temporal sequences. The benchmark automates the generation of multiple-choice QA pairs from existing video annotations, minimizing human involvement and ensuring a fair evaluation process.
  • MSVD-QA [302] The MSVD dataset is a collection of 1,970 video clips with descriptive captions, initially for video captioning. It features diverse real-world scenarios and assesses multimodal learning models’ capabilities in understanding video content and generating natural language descriptions.
  • MSRVTT-QA [302] The MSR-VTT dataset comprises 10,000 video clips with 20 human-transcribed sentences each, focusing on connecting video content with language descriptions. It evaluates multimodal learning models’ ability to comprehend video information and translate it into coherent captions, testing their video understanding and language generation skills in a more complex and diverse environment.
7.2.2 Evaluation Metric for Multimodal Datasets
The evaluation metrics for multimodal datasets include Relative Score, Accuracy, ROUGE-L, M-Avg, G-Avg, WUPS. Several common metrics, including Accuracy, ROUHE-L, have been introduced in Sec. 7.1.7. Here, we only introduce the special metrics of multimodal datasets, which include Relativa Score, M-Avg, G-Avg, WUPS as follows:
@@ 7.2.2 Evaluation Metric for Multimodal Datasets
  • Relative Score This metric is used in LLaVA-Bench to evaluate the performance of multimodal models by comparing their outputs to a reference model, typically text-based GPT-4. It is calculated as the percentage ratio of the candidate model’s score to the reference model’s score, based on dimensions such as helpfulness, relevance, accuracy, and level of detail.
@@
  • M-Avg Multiple-Choice Average is calculated as the mean accuracy across all multiple-choice tasks in the MLVU benchmark. The accuracy for each task is determined by the proportion of correctly predicted answers compared to the total number of questions within that task.
@@
  • G-Axg Generation Average s calculated as the mean score across all generation tasks in the MLVU benchmark. Each task is evaluated on multiple dimensions (e.g., Accuracy, Relevance, Completeness, and Reliability) using GPT-4, with scores ranging from 1 to 5. The overall score for each task is the average of these dimensions, and G-Avg is the mean of these task-level scores.
@@
  • WUPS [303] Wu-Palmer Similarity measures the semantic similarity between two words based on their positions in a taxonomy (e.g., WordNet). It calculates how closely related two words are by considering their least common ancestor (LCS).
@@ 8 Conclusion
Advancements in LLMs have driven significant progress on various fields, but their high computational and memory demands during inference pose challenges, especially for long-context and real-time applications. KV cache management offers an effective solution by optimizing memory, reducing redundant computation, and improving performance. This survey reviews KV cache management strategies across token-level, model-level, and system-level optimizations. Token-level optimizations focus on fine-grained control of KV cache through selection, budget allocation, merging, quantization, and low-rank decomposition, enabling efficient resource allocation without altering model architectures. Model-level optimizations leverage architectural innovations, such as attention grouping and non-transformer designs, to enhance the efficiency of KV reuse. System-level optimizations further complement these efforts by employing advanced memory management, scheduling techniques, and hardware-aware designs to optimize resource utilization across diverse computing environments.
Despite the progress made, substantial opportunities remain for future exploration. Key areas include the development of real-time, task-specific budget allocation strategies, dynamic workload handling, advanced distributed coordination for KV cache in multi-node systems, and hardware-aware innovations to leverage emerging architectures like computational storage and processing-in-memory. Additionally, integrating reinforcement learning and adaptive algorithms could enable more intelligent and responsive KV cache management, further enhancing LLM efficiency across diverse deployment scenarios.

### v3

@@ 7.3 Evaluation Metric
  • Partial Match (PM) metric evaluates model output similarity by allowing partial credit for overlaps with the reference, unlike strict metrics like Exact Match (EM).
@@
  • Recall measures a model’s ability to retrieve all relevant instances, calculated as the ratio of correctly retrieved items to total relevant items.
@@
  • F1 combines Precision and Recall using their harmonic mean, offering a balanced evaluation by considering both false positives and negatives.
  • BLEU [302] evaluates machine translation by comparing n-gram overlap with references, penalizing short outputs for fluency.
  • SacreBLEU [303] comprehensively standardizes BLEU by precisely fixing essential preprocessing steps for reliable consistent machine translation evaluations.
  • Rouge [304] and its variants measure the performance of models by calculating the overlap between the model output and the reference answer with unigram(Rouge-1), bigram(Rouge-2), LCS(Rouge-L).
@@
  • Edit Similarity measures text sequence similarity by the minimum edits needed to transform one sequence into another, derived from edit distance concepts.
  • Pass@k [307] evaluates the performance of a model by measuring the percentage of cases in which at least one of the top $k$ generated outputs contains a correct solution.
  • Exponential Similarity is a robust metric that measures the similarity between two items by exponentially weighting their differences, giving more importance to smaller discrepancies.
@@
  • Mean Reciprocal Rank (MRR) evaluates ranked results in information retrieval by averaging the reciprocal rank of the first relevant item across queries.
@@
  • Relative Score evaluates multi-modal models in LLaVA-Bench by comparing outputs to a reference model like GPT-4, calculating the percentage ratio based on helpfulness, relevance, accuracy, and detail.
@@
  • M-Avg (Multiple-Choice Average) is the mean accuracy across all multiple-choice tasks in the MLVU benchmark, based on the proportion of correct answers.
@@
  • G-Avg (Generation Average) is the mean score of generation tasks in the MLVU benchmark, evaluated on dimensions like accuracy and relevance using GPT-4, with scores from 1 to 5.
@@
  • WUPS [308] measures semantic similarity between words based on their taxonomy positions, using the least common ancestor.
@@ 8 Conclusion
Advancements in LLMs have driven significant progress in various fields, but their high computational and memory demands during inference pose challenges, especially for long-context and real-time applications. KV cache management offers an effective solution by optimizing memory, reducing redundant computation, and improving performance. This survey reviews KV cache management strategies across token-level, model-level, and system-level optimizations. Token-level optimizations focus on fine-grained control of KV cache through selection, budget allocation, merging, quantization, and low-rank decomposition, enabling efficient resource allocation without altering model architectures. Model-level optimizations leverage architectural innovations, such as attention grouping and non-transformer designs, to enhance the efficiency of KV reuse. System-level optimizations further complement these efforts by employing advanced memory management, scheduling techniques, and hardware-aware designs to optimize resource utilization across diverse computing environments.
Future directions for KV cache management research include several key areas. First, cross-category integration of token-level, model-level, and system-level optimizations presents significant opportunities but remains underexplored due to the complexity of evaluating combinatorial configurations. Second, real-world case studies are essential to understand how KV cache techniques perform in production environments, shedding light on practical trade-offs, domain-specific priorities, and implementation challenges. Third, domain-specific optimizations, such as retaining critical tokens in healthcare or applying structure-aware strategies in legal and scientific applications, can improve efficiency by tailoring techniques to unique requirements. Fourth, privacy and security considerations, including privacy-aware eviction algorithms and secure isolation mechanisms, are critical for protecting sensitive data in multi-tenant environments. Finally, addressing implementation complexity, scalability, and integration with existing frameworks will be vital to ensure the widespread adoption and practical use of KV cache strategies across diverse applications.
Acknowledegments
We would like to thank the reviewers and editors of TMLR for their constructive comments. Prof. Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab and 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA. Prof. Qing Li is supported by the Hong Kong Research Grants Council under General Research Fund (project no. 15200023) and Research Impact Fund (project no. R1015-23). Dr. Haoyang Li is supported by research funds P0052504, P0053707, and P0057770.